Calculate the OH-1 concentration in 0.0325 M HCl.
1
Expert's answer
2020-05-01T14:42:17-0400

pH+pOH=14

pH=−log[H3O+]

pOH=−log[OH]

HCl is completely ionized in water because HCl is a strong acid so the H3O+ concentration will also be 0.0325 mol/L.

pH=−log[0.0325]

pH=1.48

1.48 + pOH = 14

pOH= 12.52

12.52 = −log[OH]

10−(12.52) = [OH]

[OH] = 10−(12.52) = 3.02×10−13 mol/L

Answer: 3.02×10−13 mol/L